The Measure of the supplement of an angle is 12 more than four times the complement. Find the measure of the angle.
Found 2 solutions by ankor@dixie-net.com, chaochoa:
Answer by ankor@dixie-net.com(12692)

The Measure of the supplement of an angle is 12 more than four times the complement. Find the measure of the angle.
:
Let x = an angle
The supplement = (180-x
The complement = (90-x)
:
write what it says:
"supplement of an angle is 12 more than four times the complement.
(180 - x) = 4(90 - x) + 12
180 - x = 360 - 4x + 12
180 - x = 372 - 4x
4x - x = 372 - 180
3x = 192
x = 192/3
x = 64 degrees is the angle
:
:
Check solution:
Supplement: 180 - 64 = 116
Complement: 90 - 64 = 26
:
Using the statement:
"supplement of an angle is 12 more than four times the complement."
116 = 4(26) + 12
116 = 104 + 12, confirms our solution

Answer by chaochoa(1)

(180-x)=4(90-x)+12
180-x=360-4x+12
-x+4x=360-180+12
3x=180+12
3x=192
x=64